OpenStudy (anonymous):

I don't like these things: Identify the equation for the hyperbola with vertices at (3, 0) and (-3, 0), and an asymptote at y = 7/3x. 7x^2 - 49y^2 = 441 7x^2 - 9y^2 = 441 49x^2 - 9y^2 = 441 49x^2 + 9y^2 = 441

OpenStudy (doc.brown):

Here are my notes on hyperbolas Hyperbola\(\dfrac{x^2}{a^2}-\dfrac{y^2}{b^2}=1\) Vertices \((\pm a,0)\) Asymptotes \(y=\dfrac{\pm b}{a}x\) Eccentricity \(e=\dfrac{\sqrt{a^2+b^2}}{a}\) Length Latus Rectum = \(\dfrac{2b^2}{a}\) Focal Length \(f=\sqrt{a^2+b^2}\)